# Generalizable, Fast, and Accurate DeepQSPR with fastprop

### v3

@@ Acknowledgements
pre 3. Estrada E, Torres L, Rodriguez L, Gutman I (1998) An atom-bond connectivity index: Modelling the enthalpy of formation of alkanes
@@
pre 6. Yang K, Swanson K, Jin W, et al (2019) Analyzing learned molecular representations for property prediction. Journal of Chemical Information and Modeling 59:3370–3388. https://doi.org/10.1021/acs.jcim.9b00237
@@
pre 7. Heid E, Greenman KP, Chung Y, et al (2024) Chemprop: A machine learning package for chemical property prediction. Journal of Chemical Information and Modeling 64:9–17. https://doi.org/10.1021/acs.jcim.3c01250
@@
pre 8. Song Y, Zheng S, Niu Z, et al (2021) Communicative representation learning on attributed molecular graphs. In: Proceedings of the twenty-ninth international joint conference on artificial intelligence. Yokohama, Yokohama, Japan
@@
pre 9. Zhou G, Gao Z, Ding Q, et al (2023) Uni-mol: A universal 3D molecular representation learning framework. In: The eleventh international conference on learning representations
@@
pre 10. Chen J, Schwaller P (2023) Molecular hypergraph neural networks
@@
pre 11. Zhao B, Xu W, Guan J, Zhou S (2023) Molecular property prediction based on graph structure learning. arXiv preprint arXiv:231216855
@@
pre 12. Wang Z, Jiang T, Wang J, Xuan Q (2024) Multi-modal representation learning for molecular property prediction: Sequence, graph, geometry
@@
pre 13. Zhu Y, Chen D, Du Y, et al (2024) Molecular contrastive pretraining with collaborative featurizations. Journal of Chemical Information and Modeling 64:1112–1122. https://doi.org/10.1021/acs.jcim.3c01468
@@
pre 14. Schaduangrat N, Anuwongcharoen N, Charoenkwan P, Shoombuatong W (2023) DeepAR: A novel deep learning-based hybrid framework for the interpretable prediction of androgen receptor antagonists. Journal of Cheminformatics 15:50. https://doi.org/10.1186/s13321-023-00721-z
@@
pre 15. Tilborg D van, Brinkmann H, Criscuolo E, et al (2024) Deep learning for low-data drug discovery: Hurdles and opportunities. ChemRxiv. https://doi.org/10.26434/chemrxiv-2024-w0wvl
pre 16. McGibbon M, Shave S, Dong J, et al (2023) From intuition to AI: Evolution of small molecule representations in drug discovery. Briefings in Bioinformatics 25: https://doi.org/10.1093/bib/bbad422
@@
pre 17. Moriwaki H, Tian Y-S, Kawashita N, Takagi T (2018) Mordred: A molecular descriptor calculator. Journal of Cheminformatics 10: https://doi.org/10.1186/s13321-018-0258-y
pre 18. Falcon W, The PyTorch Lightning team (2019) PyTorch Lightning
@@
pre 19. Weininger D (1988) SMILES, a chemical language and information system. 1. Introduction to methodology and encoding rules. Journal of Chemical Information and Computer Sciences 28:31–36. https://doi.org/10.1021/ci00057a005
@@
pre 20. Comesana AE, Huntington TT, Scown CD, et al (2022) A systematic method for selecting molecular descriptors as features when training models for predicting physiochemical properties. Fuel 321:123836. https://doi.org/https://doi.org/10.1016/j.fuel.2022.123836
@@
pre 21. Wu J, Wang S, Zhou L, et al (2020) Deep-learning architecture in QSPR modeling for the prediction of energy conversion efficiency of solar cells. Industrial & Engineering Chemistry Research 59:18991–19000. https://doi.org/10.1021/acs.iecr.0c03880
@@
pre 22. Esaki T, Ohashi R, Watanabe R, et al (2019) Computational model to predict the fraction of unbound drug in the brain. Journal of Chemical Information and Modeling 59:3251–3261. https://doi.org/10.1021/acs.jcim.9b00180
@@
pre 23. Yalamanchi KK, Kommalapati S, Pal P, et al (2023) Uncertainty quantification of a deep learning fuel property prediction model. Applications in Energy and Combustion Science 16:100211. https://doi.org/https://doi.org/10.1016/j.jaecs.2023.100211
@@
pre 24. Wu Z, Ramsundar B, Feinberg EN, et al (2018) MoleculeNet: A benchmark for molecular machine learning
@@
pre 25. Ramakrishnan R, Hartmann M, Tapavicza E, Lilienfeld OA von (2015) Electronic spectra from TDDFT and machine learning in chemical space. The Journal of Chemical Physics 143: https://doi.org/10.1063/1.4928757
@@
pre 26. Ramakrishnan R, Dral PO, Rupp M, Lilienfeld OA von (2014) Quantum chemistry structures and properties of 134 kilo molecules. Scientific Data 1: https://doi.org/10.1038/sdata.2014.22
@@
pre 27. Burns JW, Spiekermann KA, Bhattacharjee H, et al (2023) Machine learning validation via rational dataset sampling with astartes. Journal of Open Source Software 8:5996. https://doi.org/10.21105/joss.05996
@@
pre 28. Notwell JH, Wood MW (2023) ADMET property prediction through combinations of molecular fingerprints
@@
pre 29. Schaduangrat N, Anuwongcharoen N, Charoenkwan P, Shoombuatong W (2023) DeepAR: A novel deep learning-based hybrid framework for the interpretable prediction of androgen receptor antagonists. Journal of Cheminformatics 15: https://doi.org/10.1186/s13321-023-00721-z
@@
pre 30. Saldana DA, Starck L, Mougin P, et al (2011) Flash point and cetane number predictions for fuel compounds using quantitative structure property relationship (QSPR) methods. Energy & Fuels 25:3900–3908. https://doi.org/10.1021/ef200795j
@@
pre 31. Das DD, St. John PC, McEnally CS, et al (2018) Measuring and predicting sooting tendencies of oxygenates, alkanes, alkenes, cycloalkanes, and aromatics on a unified scale. Combustion and Flame 190:349–364. https://doi.org/10.1016/j.combustflame.2017.12.005
@@
pre 32. Broccatelli F, Carosati E, Neri A, et al (2011) A novel approach for predicting p-glycoprotein (ABCB1) inhibition using molecular interaction fields. Journal of Medicinal Chemistry 54:1740–1751. https://doi.org/10.1021/jm101421d
@@
pre 33. Huang K, Fu T, Gao W, et al (2021) Therapeutics data commons: Machine learning datasets and tasks for therapeutics. CoRR abs/2102.09548: https://doi.org/10.48550/arXiv.2102.09548
@@
pre 34. Burns JW, Rogers DM (2023) QuantumScents: Quantum-mechanical properties for 3.5k olfactory molecules. Journal of Chemical Information and Modeling 63:7330–7337. https://doi.org/10.1021/acs.jcim.3c01338
@@
pre 35. Szymański P, Kajdanowicz T (2017) A network perspective on stratification of multi-label data. In: Luís Torgo PB, Moniz N (eds) Proceedings of the first international workshop on learning with imbalanced domains: Theory and applications. PMLR, pp 22–35
@@
pre 36. Carbonell P, Carlsson L, Faulon J-L (2013) Stereo signature molecular descriptor. Journal of Chemical Information and Modeling 53:887–897. https://doi.org/10.1021/ci300584r
@@
pre 37. Lundberg S, Lee S-I (2017) A unified approach to interpreting model predictions

### v4

@@ Acknowledgements
pre 3. Estrada’ E, Torres’ L, Rodriguez’ L (1998) An atom-bond connectivity index: Modelling the.enthalpy of formation of alkanes. Indian Journal of Chemistry 37:849–855
@@
pre 6. Coley CW, Barzilay R, Jaakkola TS, et al (2017) Prediction of organic reaction outcomes using machine learning. ACS Central Science 3:434–443. https://doi.org/10.1021/acscentsci.7b00064
@@
pre 7. Yang K, Swanson K, Jin W, et al (2019) Analyzing learned molecular representations for property prediction. Journal of Chemical Information and Modeling 59:3370–3388. https://doi.org/10.1021/acs.jcim.9b00237
@@
pre 8. Heid E, Greenman KP, Chung Y, et al (2024) Chemprop: A machine learning package for chemical property prediction. Journal of Chemical Information and Modeling 64:9–17. https://doi.org/10.1021/acs.jcim.3c01250
@@
pre 9. Song Y, Zheng S, Niu Z, et al (2021) Communicative representation learning on attributed molecular graphs. In: Proceedings of the twenty-ninth international joint conference on artificial intelligence. Yokohama, Yokohama, Japan
@@
pre 10. Zhou G, Gao Z, Ding Q, et al (2023) Uni-mol: A universal 3D molecular representation learning framework. In: The eleventh international conference on learning representations
@@
pre 11. Chen J, Schwaller P (2023) Molecular hypergraph neural networks. https://doi.org/10.48550/arXiv.2312.13136
@@
pre 12. Zhao B, Xu W, Guan J, Zhou S (2023) Molecular property prediction based on graph structure learning. arXiv preprint arXiv:231216855
@@
pre 13. Wang Z, Jiang T, Wang J, Xuan Q (2024) Multi-modal representation learning for molecular property prediction: Sequence, graph, geometry. https://doi.org/10.48550/ARXIV.2401.03369
@@
pre 14. Zhu Y, Chen D, Du Y, et al (2024) Molecular contrastive pretraining with collaborative featurizations. Journal of Chemical Information and Modeling 64:1112–1122. https://doi.org/10.1021/acs.jcim.3c01468
@@
pre 15. Schaduangrat N, Anuwongcharoen N, Charoenkwan P, Shoombuatong W (2023) DeepAR: A novel deep learning-based hybrid framework for the interpretable prediction of androgen receptor antagonists. Journal of Cheminformatics 15:50. https://doi.org/10.1186/s13321-023-00721-z
@@
pre 16. Tilborg D van, Brinkmann H, Criscuolo E, et al (2024) Deep learning for low-data drug discovery: Hurdles and opportunities. ChemRxiv. https://doi.org/10.26434/chemrxiv-2024-w0wvl
@@
pre 17. Moriwaki H, Tian Y-S, Kawashita N, Takagi T (2018) Mordred: A molecular descriptor calculator. Journal of Cheminformatics 10:4. https://doi.org/10.1186/s13321-018-0258-y
pre 18. McGibbon M, Shave S, Dong J, et al (2023) From intuition to AI: evolution of small molecule representations in drug discovery. Briefings in Bioinformatics 25:bbad422. https://doi.org/10.1093/bib/bbad422
pre 19. Falcon W, The PyTorch Lightning team (2019) PyTorch Lightning
@@
pre 20. Weininger D (1988) SMILES, a chemical language and information system. 1. Introduction to methodology and encoding rules. Journal of Chemical Information and Computer Sciences 28:31–36. https://doi.org/10.1021/ci00057a005
@@
pre 21. Comesana AE, Huntington TT, Scown CD, et al (2022) A systematic method for selecting molecular descriptors as features when training models for predicting physiochemical properties. Fuel 321:123836. https://doi.org/https://doi.org/10.1016/j.fuel.2022.123836
@@
pre 22. Wu J, Wang S, Zhou L, et al (2020) Deep-learning architecture in QSPR modeling for the prediction of energy conversion efficiency of solar cells. Industrial and Engineering Chemistry Research 59:18991–19000. https://doi.org/10.1021/acs.iecr.0c03880
@@
pre 23. Esaki T, Ohashi R, Watanabe R, et al (2019) Computational model to predict the fraction of unbound drug in the brain. Journal of Chemical Information and Modeling 59:3251–3261. https://doi.org/10.1021/acs.jcim.9b00180
@@
pre 24. Yalamanchi KK, Kommalapati S, Pal P, et al (2023) Uncertainty quantification of a deep learning fuel property prediction model. Applications in Energy and Combustion Science 16:100211. https://doi.org/https://doi.org/10.1016/j.jaecs.2023.100211
@@
pre 25. Wu Z, Ramsundar B, Feinberg EN, et al (2018) MoleculeNet: A benchmark for molecular machine learning. https://doi.org/10.48550/arXiv.1703.00564
@@
pre 26. Ramakrishnan R, Hartmann M, Tapavicza E, von Liliendfeld O (2015) Electronic spectra from TDDFT and machine learning in chemical space. The Journal of Chemical Physics 143: https://doi.org/10.1063/1.4928757
@@
pre 27. Ramakrishnan R, Dral P, Rupp M, von Liliendfeld O (2014) Quantum chemistry structures and properties of 134 kilo molecules. Scientific Data 1: https://doi.org/10.1038/sdata.2014.22
@@
pre 28. Burns J, Spiekermann K, Bhattacharjee H, et al (2023) Machine learning validation via rational dataset sampling with astartes. Journal of Open Source Software 8:5996. https://doi.org/10.21105/joss.05996
@@
pre 29. Notwell JH, Wood MW (2023) ADMET property prediction through combinations of molecular fingerprints. https://doi.org/10.48550/arXiv.2310.00174
@@
pre 30. Schaduangrat N, Anuwongcharoen N, Charoenkwan P, Shoombuatong W (2023) DeepAR: A novel deep learning-based hybrid framework for the interpretable prediction of androgen receptor antagonists. Journal of Cheminformatics 15:50. https://doi.org/10.1186/s13321-023-00721-z
@@
pre 31. Saldana DA, Starck L, Mougin P, et al (2011) Flash point and cetane number predictions for fuel compounds using quantitative structure property relationship (QSPR) methods. Energy & Fuels 25:3900–3908. https://doi.org/10.1021/ef200795j
@@
pre 32. Das DD, St. John PC, McEnally CS, et al (2018) Measuring and predicting sooting tendencies of oxygenates, alkanes, alkenes, cycloalkanes, and aromatics on a unified scale. Combustion and Flame 190:349–364. https://doi.org/10.1016/j.combustflame.2017.12.005
@@
pre 33. Broccatelli F, Carosati E, Neri A, et al (2011) A novel approach for predicting p-glycoprotein (ABCB1) inhibition using molecular interaction fields. Journal of Medicinal Chemistry 54:1740–1751. https://doi.org/10.1021/jm101421d
@@
pre 34. Huang K, Fu T, Gao W, et al (2021) Therapeutics data commons: Machine learning datasets and tasks for therapeutics. CoRR abs/2102.09548: https://doi.org/10.48550/arXiv.2102.09548
@@
pre 35. Burns JW, Rogers DM (2023) QuantumScents: Quantum-mechanical properties for 3.5k olfactory molecules. Journal of Chemical Information and Modeling 63:7330–7337. https://doi.org/10.1021/acs.jcim.3c01338
@@
pre 36. Szymański P, Kajdanowicz T (2017) A network perspective on stratification of multi-label data. In: Luís Torgo PB, Moniz N (eds) Proceedings of the first international workshop on learning with imbalanced domains: Theory and applications. PMLR, pp 22–35
@@
pre 37. Carbonell P, Carlsson L, Faulon J-L (2013) Stereo signature molecular descriptor. Journal of Chemical Information and Modeling 53:887–897. https://doi.org/10.1021/ci300584r
@@
pre 38. Lundberg S, Lee S-I (2017) A unified approach to interpreting model predictions. https://doi.org/10.48550/ARXIV.1705.07874